# Using analytic models to describe effective PDFs

### v1

@@ V.1 Run time
We start by presenting, in Tab. (1), the run-time difference when performing a call to our PDFs and LHAPDF, for sets of points in the $(x,Q^{2})$ space. As we can appreciate, for $N_{points}\leq 10^{3-4}$ the gain is quite substantial. This is due to the fact that calling LHAPDF has an overhead time for loading and reading the grid, which albeit small, dominates the total run time when one evaluates a small number of points. Therefore, if we were interested in evaluating a handful of points we would be wise in choosing our ML-PDFs over LHAPDF. As we move to larger number of points $N_{points}\approx 10^{5}$ the weight of this overhead starts to dilute, and once we pass the mark $N_{points}\approx 10^{8}$, we reach a region where the run time of both methods seem to be comparable. As a consequence, it appears that for a code requiring Monte-Carlo integration, it would not make any difference to apply one or the other. We further investigate this in the next sub-section.
@@
$10^{3}$ $4.81\cdot 10^{-2}$ $5.78\cdot 10^{-4}$ 98.8
$10^{4}$ $4.60\cdot 10^{-2}$ $5.26\cdot 10^{-3}$ 88.6
$10^{5}$ $9.14\cdot 10^{-2}$ $5.27\cdot 10^{-2}$ 42.3
$10^{6}$ $0.57$ $0.52$ 7.5
$10^{7}$ $5.29$ $5.25$ 1
$10^{8}$ $52.49$ $52.43$ 0.1
@@
V.2 Impact in physical observables
@@
We start by considering pion production in unpolarised proton-proton collisions at $\sqrt{S_{\rm c.m.}}=7$ TeV in the central rapidity region ($|\eta^{\pi}|<0.5$). These simulations make use of a sequential code based on Ref. Jager et al. (2003), modified to use the needed PDFs and fragmentation functions (FFs). The Monte-Carlo integrator is VEGAS, with $10^{5}$ points, the FF set used is DEHSS2014 Hernández-Pinto et al. (2016), and all factorization and renormalization scales have been taken to be equal to the transverse momentum of the pion (denoted $p_{T}$). In the upper plot of Fig. 8, we present the $p_{T}$ spectrum using our ML-PDFs (black dots) and HERAPDF2.0 (red line), while the lower plot depicts the ratio between these two quantities. The (light-blue) uncertainty band in the lower plot comes from the Hessian set of theoretical uncertainties of HERAPDF2.0; it is also present (but hardly noticeable) in the upper plot. From the lower panel, we can appreciate that the difference between both simulations is well below $2\,\%$, and mostly stays within the PDF uncertainty bands. This is much smaller than the typical perturbative error of NLO QCD calculations (generally larger than ${\cal O}(20\,\%)$). Regarding the CPU time, it took 10 minutes 28 seconds using LHAPDF and 10 minutes 3 seconds with our ML-PDFs, running on the same system described before. From Tab. 1, and considering the number of iterations used, one would expect a much larger gain. The outcome of this run serves to emphasize that the matter is far more subtle. It is worth highlighting that this reduction in the computational time is only due to our optimized ML-PDF, since all the other ingredients of the calculation remain the same in both simulations.
@@
In this case, we chose to run $10^{8}$ points both in order to be in what we expect (from Tab. 1) to be a very low time-gain region and because the observable requires a large number of points to produce physical results. We show the outcome in Fig. 9. As before, the upper plot presents the cross-section, while the lower plot demonstrates the ratio between the two computations. Given the $p_{T}$ range explored, there are some oscillations of the output associated to statistical fluctuations. In fact, using $10^{7}$ points gave unphysical results for some $p_{T}$ values. We can then say that this observable would greatly benefit from using a higher number of points. For our purpose, $10^{8}$ points is enough. In the $p_{T}\leq 10$ GeV region, that dominates the cross-section, the differences are within $5\%$ and are feasible of further reduction. The runtime for LHAPDF was of 461 minutes, almost 7 hours, while the ML-PDF run took 403 minutes, almost a full hour less. We summarise the runtime of the cross-sections in Tab. 2. This gain is much larger than in the previous test, and far more than expected from the table. Given the duration of this run, we refrained from including the PDF theoretical uncertainties.
Observable LHAPDF (s)   ML-PDFs (s)
$p+p\to\pi$ $628.320$ $603.552$
$p+p\to\gamma+\pi$ $27648.795$ $24179.811$
@@ VI Conclusions
In this work, we provide an analytic approximation to PDFs using ML-techniques to adjust both their $x$ and $Q$-dependence. Our starting hypothesis was the assumption that the $x$-dependence could be reproduced by Eulerian-like functions, whilst all the $Q$-dependence is embodied within the coefficients.
By doing so, we obtain a reliable approximation to HERAPDF2.0, taking into consideration up to NLO QCD corrections. We show that the integral error is under control for an ample range of $x$ and $Q$ values. In fact, our ML-PDFs were tested for $Q\in[Q_{0},1000\,{\rm GeV}]$, presenting (for most of the distributions) deviations w.r.t. HERAPDF2.0 below percent level. It is important to highlight that this is comparable with the error of the PDF sets themselves, and far smaller than the uncertainties introduced by truncating the perturbative expansion of observables in QCD. In fact, we find that the error induced by the ML-PDF in $p+p\to\pi$ is ${\cal O}(1\,\%)$ whilst for $p+p\to\gamma+\pi$ is ${\cal O}(5\,\%)$  [^2] In this work, we neglect to study the uncertainty in the parameters of the ML-PDFs. Such a study is beyond the scope of the present work, and we leave it to future studies.. These errors are much smaller than the ${\cal O}(20-50\,\%)$or even larger- uncertainties arising from scale variations. Furthermore, we found a non-negligible reduction of the runtime: ${\cal O}(4\,\%)$ for $p+p\to\pi$ and ${\cal O}(13\,\%)$ for $p+p\to\gamma+\pi$.
To conclude, we want to emphasize that the strategy explained in this article is fully applicable to any PDF set, and even to FFs, at any perturbative order (NLO, NNLO and beyond), due to the fact that we made no assumptions about the order at which DGLAP evolution is truncated. Our only hypothesis is that the $Q$-dependence is fully embodied within the coefficients $\{A,B,\ldots\}$ described in Sec. IV. Following this approach, we can by-pass DGLAP evolution and speed up the calculation of PDFs (and, eventually, FFs). This does not only imply a reduction of the CPU time, but also leads to a reduction of the CO2 footprint and points towards more efficient and sustainable practises in high-energy physics (HEP) Banerjee et al. (2023). For this reason, further developments to obtain analytic approximations to PDFs/FFs are expected and this work can be considered as a first proof-of-concept paving the way for future improvements. As final remark, we want to stress that our approach does not replace the PDF extractions from data; in fact, they constitute an essential input and the basis for ML-PDF determinations as the one presented here.

### v2

@@ V.1 Run time
We start by presenting, in Tab. (1), the run-time difference when performing a call to our PDFs and LHAPDF, for sets of points in the $(x,Q^{2})$ space. As we can appreciate, for $N_{points}\leq 10^{3-4}$ the gain is quite substantial. This is due to the fact that calling LHAPDF has an overhead time for loading and reading the grid, which albeit small, dominates the total run time when one evaluates a small number of points. Therefore, if we were interested in evaluating a handful of points we would be wise in choosing our ML-PDFs over LHAPDF. As we move to larger number of points $N_{points}\approx 10^{5}$ the weight of this overhead starts to dilute, and once we pass the mark $N_{points}\approx 10^{6}$, we reach a region where we are essentially comparing only the execution time of interpolation versus evaluation. In the last column we display the percentual time gain, defined as
@@
A negative gain would mean that the interpolation is faster than the direct evaluation. As can be seen from the table, the gain seems to approach a plateau around $50\%$ (i.e. our ML-PDF are two times faster than LHAPDF), which is quite sizable.
[…]
$10^{3}$ $3.76\cdot 10^{-2}$ $2.92\cdot 10^{-4}$ 99.22
$10^{4}$ $4.20\cdot 10^{-2}$ $2.50\cdot 10^{-3}$ 94.05
$10^{5}$ $8.94\cdot 10^{-2}$ $2.50\cdot 10^{-2}$ 72.10
$10^{6}$ $0.56$ $0.25$ 55.46
$10^{7}$ $5.25$ $2.50$ 52.49
$10^{8}$ $52.04$ $24.92$ 52.11
@@
From these numbers it appears that, for a code requiring Monte-Carlo integration, it would be beneficial to use something akin to our ML-PDFs. We further investigate this in the next sub-section.
V.2 Validity of the sum rules
One important physically-motivated cross-check of our analytic ML-PDFs are the sum rules. For any energy scale $Q$, the sum rules are given by
[…]
for the up and down valence quarks, and
[…]
for all other quarks. Also, the PDFs must fulfill that the total momenta carried by the constituents partons equal the momentum of the hadron. This is encoded in the momentum sum rule
[…]
where the sum is carried out over all the active flavours of quarks. These constraints are often imposed when performing the PDF extraction from fits to experimental data. In the case of HERAPDF2.0, Eq. (44) is automatically fulfilled by construction (no $q-\bar{q}$ distinction for sea quarks), and this also holds for our analytic ML-PDFs.
Figure 9: Sum rules for up (upper) and down (lower) valence distributions as a function of $Q$. We show the values calculated with HERAPDF2.0 (blue) and our analytic ML-PDF (orange), and also the theoretical exact value (green).
Figure 10: Total momentum conservation rule, as a function of $Q$. We show the values calculated with HERAPDF2.0 (blue) and our analytic ML-PDF (orange), and also the theoretical exact value (green).
In Fig. 9 we compare the result of Eqs. (42) and (43) for HERAPDF2.0 (blue) and our analytic ML-PDFs (orange) w.r.t. the theoretical value (green line). In both cases, we consider $Q\in[10\,{\rm GeV},1000\,{\rm GeV}]$ and we find that our analytic ML-PDF leads to results closer to the theoretical value. In concrete, for up valence quark, we get deviations of ${\cal O}(0.1\,\%)$ for ML-PDFs and ${\cal O}(0.2\,\%)$ for HERAPDF2.0, showing the last one a trend to depart from the theoretical value in the high $Q$ region. Similarly, for down valence distributions, the predictions obtained with ML-PDFs oscillate around 1 within a band of ${\cal O}(0.2\,\%)$, while HERAPDF2.0 deviates more than ${\cal O}(0.3\,\%)$ for higher $Q$-values.
Finally, in Fig. 10, we show the results of computing Eq. (45) with our analytic ML-PDFs (orange) and HERAPDF2.0 (blue). In this case, HERAPDF2.0 approaches better to the theoretical value, with discrepancies smaller than ${\cal O}(0.5\,\%)$. The predictions obtained with our analytic ML-PDFs shows deviations of ${\cal O}(0.6\,\%)$ below 60 GeV, reaching up to ${\cal O}(1.5\,\%)$ for $Q\approx 500$ GeV. This behaviour is driven by gluon PDF, since this parton carries more than $45\,\%$ of proton momentum and our fit show an increasing error for large $Q$ (see Figs. 7 and 8). In any case, as we emphasized in the Introduction, this constitutes a first proof-of-concept and we can claim that our ML-PDFs successfully passed all the cross-check with a percent (and even sub-percent) level precision.
V.3 Impact in physical observables
@@
We start by considering pion production in unpolarised proton-proton collisions at $\sqrt{S_{\rm c.m.}}=7$ TeV in the central rapidity region ($|\eta^{\pi}|<0.5$). These simulations make use of a sequential code based on Ref. Jager et al. (2003), modified to use the needed PDFs and fragmentation functions (FFs). The Monte-Carlo integrator is VEGAS, with $10^{5}$ points, the FF set used is DEHSS2014 Hernández-Pinto et al. (2016), and all factorization and renormalization scales have been taken to be equal to the transverse momentum of the pion (denoted $p_{T}$). In the upper plot of Fig. 11, we present the $p_{T}$ spectrum using our ML-PDFs (black dots) and HERAPDF2.0 (red line), while the lower plot depicts the ratio between these two quantities. The (light-blue) uncertainty band in the lower plot comes from the Hessian set of theoretical uncertainties of HERAPDF2.0; it is also present (but hardly noticeable) in the upper plot. From the lower panel, we can appreciate that the difference between both simulations is well below $2\,\%$, and mostly stays within the PDF uncertainty bands. This is much smaller than the typical perturbative error of NLO QCD calculations (generally larger than ${\cal O}(20\,\%)$). Regarding the CPU time, it took 10 minutes 28 seconds using LHAPDF and 9 minutes 19 seconds with our ML-PDFs, running on the same system described before. From Tab. 1, and considering the number of iterations used, one would expect a much larger gain. The outcome of this run serves to emphasize that the matter is far more subtle. It is worth highlighting that this reduction in the computational time is only due to our optimized ML-PDF, since all the other ingredients of the calculation remain the same in both simulations.
@@
In this case, we chose to run $10^{8}$ points both in order to be in what we expect (from Tab. 1) to be a very low time-gain region and because the observable requires a large number of points to produce physical results. We show the outcome in Fig. 12. As before, the upper plot presents the cross-section, while the lower plot demonstrates the ratio between the two computations. Given the $p_{T}$ range explored, there are some oscillations of the output associated to statistical fluctuations. In fact, using $10^{7}$ points gave unphysical results for some $p_{T}$ values. We can then say that this observable would greatly benefit from using a higher number of points. For our purpose, $10^{8}$ points is enough. In the $p_{T}\leq 10$ GeV region, that dominates the cross-section, the differences are within $5\%$ and are feasible of further reduction. The runtime for LHAPDF was of 207.5 minutes, almost 3.45 hours, while the ML-PDF run took 144.5 minutes, more than a full hour less. We summarise the runtime of the cross-sections in Tab. 2. This gain is much larger than in the previous test, and far more than expected from the table. Also, we notice that the predictions obtained with our ML-PDFs are compatible with the propagation of errors of HERAPDF2.0 into the observable (light blue band).
[…]
$p+p\to\pi$ $628.320$ $558.854$ $11.06$
$p+p\to\gamma+\pi$ $12452.273$ $8671.827$ $30.36$
@@ VI Conclusions
In this work, we provide an analytic approximation to PDFs using machine-assisted techniques to adjust both their $x$ and $Q$-dependence. Our starting hypothesis was the assumption that the $x$-dependence could be reproduced by Eulerian-like functions, whilst all the $Q$-dependence is embodied within the coefficients.
By doing so, we obtain a reliable approximation to HERAPDF2.0, taking into consideration up to NLO QCD corrections. We show that the integral error is under control for an ample range of $x$ and $Q$ values. In fact, our ML-PDFs were tested for $Q\in[Q_{0},1000\,{\rm GeV}]$, presenting (for most of the distributions) deviations w.r.t. HERAPDF2.0 below percent level. It is important to highlight that this is comparable with the error of the PDF sets themselves, and far smaller than the uncertainties introduced by truncating the perturbative expansion of observables in QCD. In fact, we find that the error induced by the ML-PDF in $p+p\to\pi$ is ${\cal O}(1\,\%)$ whilst for $p+p\to\gamma+\pi$ is ${\cal O}(5\,\%)$  [^2] In this work, we neglect to study the uncertainty in the parameters of the ML-PDFs. Such a study is beyond the scope of the present work, and we leave it to future studies.. These errors are much smaller than the ${\cal O}(20-50\,\%)$or even larger- uncertainties arising from scale variations. Furthermore, we found a non-negligible reduction of the runtime: ${\cal O}(11\,\%)$ for $p+p\to\pi$ and ${\cal O}(30\,\%)$ for $p+p\to\gamma+\pi$.
To conclude, we want to emphasize that the strategy explained in this article is fully applicable to any PDF set, and even to FFs, at any perturbative order (NLO, NNLO and beyond), due to the fact that we made no assumptions about the order at which DGLAP evolution is truncated. The only required ingredient is a previously existent PDF, extracted from data, to be used as input in our formalism. Our only hypothesis is that the $Q$-dependence is fully embodied within the coefficients $\{A,B,\ldots\}$ described in Sec. IV. Following this approach, we can avoid using interpolation methods to evaluate the distributions and speed up the calculation of PDFs (and, eventually, FFs), because we obtain a fit that already embodied the evolution in an analytically approximated way. This does not only bring a reduction of the CPU time, but also leads to a reduction of the CO2 footprint and points towards more efficient and sustainable practises in high-energy physics (HEP) Banerjee et al. (2023). For this reason, further developments to obtain analytic approximations to PDFs/FFs are expected and this work can be considered as a first proof-of-concept paving the way for future improvements. As final remark, we want to stress that our approach does not replace the PDF extractions from data; in fact, they constitute an essential input and the basis for ML-PDF determinations as the one presented here.